What happens to the magnitude of a the gravitational field strength if the distance from a planet doubles and the test mass quadruples?


1
Expert's answer
2021-01-07T10:55:16-0500
"g=\frac{GM}{R^2}\\g'=\frac{GM}{(2R)^2}=\frac{GM}{4R^2}=\frac{g}{4}"